OpenStudy (anonymous):

James works as a waiter.Bills where $23.59,$40.65,$30.50 And $15.68 His total tips was $17.67. What percent tip did he recieve for serving these meals. Round to nearest percent.

OpenStudy (anonymous):

=17.67/(23.59+40.65+20.50+15.68) =17.67/100.42 =.176==17.6%